# The Hidden Link Between RLHF and Contrastive Learning

### v1

@@ 3.1 Why should JSD be used instead of MINE in Alignment?
As our primary objective is to maximize MI rather than precisely estimate its value [23], we can rely on non-KL divergences, which offer favorable trade-offs. For instance, the Jensen-Shannon MI estimator [32]—similar to the binary cross-entropy used in minimizing total correlation [33]—is well-understood in neural network optimization and is more stable in practice compared to the DV-based objective [23]:
@@
The left panel of Figure 2 compares the MINE and JSD estimators with a single negative sample ($K=1$). While both estimators qualitatively capture the ground-truth MI trend, JSD does not consistently outperform MINE in terms of accuracy. However, when gradient stability is prioritized over precise MI estimation, JSD demonstrates a clear advantage [34]. To further investigate this, we analyze the gradient variance induced by each estimator, with the right panel of Figure 2 showing the gradient variance computed from the first-layer weights of a three-layer MLP trained for each estimator’s objective. Under conditions with limited negative samples, JSD exhibits significantly greater stability than MINE, making it particularly suitable for training objectives, such as Alignment, that rely on contrastive MI estimates with sparse negative sampling [12].
3.2 From DPO to Mutual-Information Optimisation (MIO)
@@
By replacing the DV/MINE [22] surrogate in Section 2 with the JS-based surrogate given in Equations (11) and (4), and substituting  (12) into  (11), we obtain the following empirical objective:
@@
Setting $M=N=1$ and constraining the critic to $T_{\phi}(\pi,\pi_{\theta})=\log\frac{\pi_{\theta}}{\pi}$ gives a closed-form loss, which we call Mutual-Information Optimisation (MIO):
@@
Why MIO is Preferable to DPO Under Low $K$.
With $K=1$ negative sample, the DV/MINE surrogate induces unstable gradients due to its high-variance denominator, whereas the JSD surrogate reduces gradient variance without compromising performance (Fig. 2, right panel). Since DPO is based on DV/MINE [22] and MIO is based on JSD [23], the policy update driven by (14) is more stable, particularly when negative comparisons per prompt are limited [12]. When few negative samples are available, the gradients from MINE become unstable, with the choice of negatives significantly affecting the estimate and amplifying gradient variance. Given that our primary goal is to obtain reliable gradients rather than precise MI estimates, the JSD estimator is preferred [34]. Due to its substantially lower variance, MIO results in smoother optimization trajectories and mitigates gradient oscillations caused by noisy negative sampling.Empirical results in Figure 5 confirm that MIO strengthens alignment with the chosen responses while reducing consistency with the rejected responses. In contrast, DPO [35] decreases consistency with both the chosen and rejected responses, which may impair the model’s reasoning and mathematical performance [17, 16].
4 Experiment
4.1 DPO Breaks When Rejection near 0, But MIO Not!
Following [20], the toy model setup, shown in Figure 3, consists of a discrete space with 4 prompts and 10 responses. The policy $\pi_{\theta}$, implemented as a three-layer MLP, processes a one-hot input vector and outputs a categorical distribution over the responses, which are divided into three categories: selected responses (first 4 dimensions), rejected responses (next 4 dimensions), and unseen responses (final 2 dimensions). Each prompt has a corresponding optimal response (e.g., response 1 is optimal for prompt 1). In DPO or MIO training, preference data is generated through a mini-batch sampling strategy, where an ideal annotator matches each prompt with its optimal response, creating a diagonal preference matrix. In each mini-batch, one additional response is randomly selected to form preference pairs, ensuring diverse gradient updates.
Figure 3: Toy model setup[20]. Left: the optimal policy where the highlighted blocks represent optimal responses. Right: preference dataset construction.
Figure 4 contrasts the learning dynamics of DPO with those of our proposed MIO objective. When the probabilities of both the chosen and rejected responses remain in a moderate range, DPO converges smoothly and the likelihood of the chosen response declines only marginally relative to that of the rejected one. Once the rejected–response probability becomes very small, however, DPO develops a pathological synchronous collapse: the probabilities of both the chosen and the rejected responses decrease in lock-step. Across all regimes we tested, MIO displays no such joint degradation. When the rejected probability is pushed towards zero, DPO and its derivatives exhibit synchronous decline of the chosen and rejected likelihoods, mirroring what is observed in practice [20, 12, 8], whereas MIO remains entirely immune to this failure mode. A formal explanation of this qualitative difference is provided in Section 5.
The toy model serves as an abstract simulation that amplifies the 3D-properties of DPO [20], which leads to a decrease in the likelihood of both chosen and rejected responses. However, MIO does not exhibit this effect. While the toy model setup differs significantly from real-world LLM training—such as in sampling frequency—it provides valuable insights. In real-world scenarios, DPO and MIO are typically trained over one epoch, with each data point used only a few times. In contrast, the toy model samples the same data points repeatedly. Conceptually, this is akin to treating each input/output as a token rather than a full prompt/response, where each token may be sampled multiple times during real-world training. The code will be provided in the supplementary materials.
4.2 Experimental Setup
We fine-tune the Mistral-7B-Base [36] on the 64 k-prompt UltraFeedback-Binarized [37] corpus, where each prompt is paired with a chosen and a rejected completion selected by GPT-4, yielding 64 k positive–negative pairs; we retain the official train/validation split of UltraFeedback-Binarized.
Model quality is assessed on eight public mathematical and reasoning suites: Hendrycks MATH [38] (12.5 k competition-level problems plus the level-5 “hard” subset), Minerva Math [39] (quantitative-reasoning items distilled from GSM8K and MATH), MultiMedQA [40] (six medical QA benchmarks spanning USMLE, PubMedQA, and consumer health queries), MathQA [41] (37 k multiple-choice word problems with executable rationales), GSM8K [42] (8.5 k grade-school arithmetic questions), AQuA-RAT [43] (100 k algebraic problems with free-text solutions), and MuSR [44] (narrative tasks requiring multistep soft reasoning over 700–1000-word stories). The code, hyperparameter settings, and evaluation metrics will be provided in the supplementary materials.
Table 1: Evaluation results on mathematical & reasoning benchmarks (the higher, the better). The best and second best numbers are marked with boldface and underline.
Method ($\downarrow$) / Benchmark ($\rightarrow$) Hendrycks Math Minerva Math Multimedqa MathQA GSM8K Aqua Rat Math Hard MUSR
SFT 0.14 8.80 49.03 29.91 32.29 19.29 2.72 39.94
DIL[4] 0.04 6.46 49.54 28.94 28.35 21.33 2.11 38.36
DPO[5] 0.04 8.14 49.21 30.72 36.47 19.69 2.72 43.65
IPO[45] 0.08 6.50 48.39 29.41 3.71 19.69 0.23 35.71
KTO[46] 0.04 10.38 49.76 30.56 41.75 21.65 3.43 43.39
NCA[12] 0.14 8.92 49.58 30.59 37.68 18.50 2.49 39.68
ORPO[47] 0.04 9.68 49.40 29.78 36.39 24.80 2.57 42.06
SimPO[6] 0.19 8.36 49.16 30.93 39.20 18.11 2.49 39.02
SLiC[48] 0.08 8.72 49.74 30.66 39.73 20.47 3.17 42.99
MIO 0.21 9.08 49.82 30.96 40.16 24.83 3.57 43.42
Across the eight mathematical and reasoning benchmarks, MIO achieves the best performance on five tasks and ranks second on two others. These results demonstrate that MIO offers superior generalization in reasoning and mathematical problem-solving compared to existing alignment methods.
4.3 MIO prevents the chosen-reward collapse
Figure 5: Training curves on Mistral-7B-Base. Unlike DPO, IPO, and ORPO, MIO increases the rewards of chosen responses while suppressing rejected ones, thereby avoiding the “synchronous collapse” observed in prior methods.
During DPO training, the rewards of all responses ($y_{w}$ and $y_{l}$) often decrease simultaneously, a harmful “unlearning” effect that degrades downstream performance [4, 6, 17]. Figure 5 shows that the proposed MIO objective completely eliminates this failure mode: the chosen-rewards increase while the rejected rewards drop slightly. Section 5 provides a theoretical explanation of this behavior.
5 Why MIO Outperforms DPO?
5.1 Failure mode of DPO
For a preference triple $(x,a^{+},a^{-})$ the DPO loss can be written as [20]:
@@ 5.1 Failure mode of DPO
$\ell_{\mathrm{DPO}}=\log\!\Bigl{(}1+\bigl{[}\alpha\,z^{\beta}\bigr{]}\Bigr{)},\qquad\alpha=\Bigl{[}\tfrac{\pi_{0}(a^{+}\!\mid x)}{\pi_{0}(a^{-}\!\mid x)}\Bigr{]}^{\beta},\;z=\tfrac{\pi_{\theta}(a^{-}\!\mid x)}{\pi_{\theta}(a^{+}\!\mid x)}.$
Denoting $\pi^{+}\!=\!\pi_{\theta}(a^{+}\!\mid x)$ and $\pi^{-}\!=\!\pi_{\theta}(a^{-}\!\mid x)$, differentiation gives
@@
Hence, we have that $\lvert\partial_{\pi^{-}}\ell\rvert/\lvert\partial_{\pi^{+}}\ell\rvert=\pi^{+}/\pi^{-}$. The gradient on the rejected probability therefore grows much faster than that on the chosen probability; as $\pi^{-}\!\to 0$ we have $\partial_{\pi^{-}}\ell\!\to\!\infty$ but $\partial_{\pi^{+}}\ell\!\to 0$. Because chosen and rejected completions usually share many tokens, the vanishing positive gradient cannot offset the dominant negative gradient, so the model is driven to lower both probabilities. This synchronous likelihood collapse—observed empirically in late-stage DPO training—degrades overall performance [17].
5.2 Stability of MIO
@@
The MIO objective for the same triple is
@@
with $\operatorname{LR}^{\pm}=\log\pi^{\pm}-\log\pi_{0}^{\pm}$. Writing $\sigma^{\pm}=\sigma\!\bigl{(}\beta\operatorname{LR}^{\pm}\bigr{)}$, one obtains
@@ 6 Conclusion
In this work, we propose a novel perspective that frames RLHF and DPO as a unified form of contrastive learning over the distributions of chosen and rejected responses, grounded in mutual information. Building on this connection, we theoretically show that RLHF does not extend the base model’s capacity but merely amplifies its existing capabilities. Furthermore, we introduce the Mutual Information Optimization (MIO), replacing the high-variance Donsker-Varadhan (DV) bound with a Jensen-Shannon (JS) bound. MIO demonstrates superior preservation of reasoning and mathematical abilities and offers greater training stability compared to existing baselines. Extensive experimental results demonstrate that MIO achieves superior performance on a wide range of mathematical and reasoning benchmarks. We hope that our work could offer useful insights for the community and help address the limitations of current RLHF methods. This work does not explore the effects of alternative mutual information estimators, which we leave as a promising direction for future research.
@@
Substitute (A.3) into (A.1).
@@
Insert (A.4) into (A.2).
@@
and using (A.3) in (A.5) we arrive at
@@
For every prompt $x$ we draw $y_{w}\sim\pi_{\text{chosen}}(\cdot\!\mid\!x)$ (positive) and $y_{l}\sim\pi_{\text{rejection}}(\cdot\!\mid\!x)$ (negative). Fix a critic $T_{\phi}\!\in\!\mathcal{F}$. With $z^{+}=T_{\phi}(x,y_{w}),\;z^{-}=T_{\phi}(x,y_{l})$, define the two DV / InfoNCE estimators:
@@
Write the logit gap $\Delta\coloneqq z^{+}-z^{-}$. Using $\sigma(u)=\tfrac{1}{1+e^{-u}}$, we have $\hat{I}^{+}=\log\sigma(\Delta),\;\hat{I}^{-}=\log\sigma(-\Delta)$.
@@
Because $\sigma(u)>0$ for all $u\in\mathbb{R}$,
@@
$\nabla_{\theta}\hat{I}^{+}\;=\;\frac{\partial\hat{I}^{+}}{\partial\Delta}\,\nabla_{\theta}\Delta,\quad\nabla_{\theta}\hat{I}^{-}\;=\;\frac{\partial\hat{I}^{-}}{\partial\Delta}\,\nabla_{\theta}\Delta.$
Combine (B.3)–(B):
@@
The scalar factor $-\sigma(\Delta)/\sigma(-\Delta)\,<0$; hence
@@
i.e., the two gradient vectors always point in opposite directions.
Consequence.
$\hat{I}^{+}(\theta_{\text{new}})>\hat{I}^{+}(\theta),\;\Longrightarrow\;\hat{I}^{-}(\theta_{\text{new}})<\hat{I}^{-}(\theta).$
Hence, maximizing the estimated mutual information between $\pi_{\theta}$ and $\pi_{\text{chosen}}$ simultaneously minimizes the estimated mutual information between $\pi_{\theta}$ and $\pi_{\text{rejection}}$. This coupling holds for every step of gradient ascent and therefore for the final optimum, completing the proof.
Appendix C Mutual-Information Estimation: Optimal Critic and Practical Surrogate

### v2

@@ 3 Methods: Mutual-Information Optimisation
As our primary objective is to maximize Mutual Information rather than precisely estimate its value (Hjelm et al., 2019), we can rely on non-KL divergences, which offer favorable trade-offs. For instance, the Jensen-Shannon MI estimator (Nowozin et al., 2016)—similar to the binary cross-entropy used in minimizing total correlation (Brakel & Bengio, 2017)—is well-understood in neural network optimization and is more stable in practice compared to the DV-based objective (Hjelm et al., 2019):
@@
By replacing the DV/MINE (Belghazi et al., 2018a) surrogate $I_{\mathrm{DV}}$ in  equation 4 with the JS-based surrogate $I_{\mathrm{JSD}}$ given in  equation 12, and substituting  equation 13 into  equation 12, we obtain the following empirical objective:
@@
Setting $M=N=1$ and constraining the critic to a log ratio family (See in Appendix  C.2) gives a closed-form loss, which we call Mutual-Information Optimisation (MIO):
@@
3.2 Failure Mode of DPO when $\pi_{\theta}(y\mid x)$ is Close to Zero
Essentially, DPO estimates the policy gradient through the DV/MINE mutual information estimator. Consequently, it suffers from the same failure mode: when $\pi_{\theta}(y\mid x)$ becomes too small, the policy gradient vanishes. This effect is particularly pronounced for the rejected sample $y_{l}$ in preference alignment, since training explicitly encourages reducing its probability, making $\pi_{\theta}(y_{l}\mid x)$ more likely to approach zero over the course of optimization.
For a preference triple $(x,y_{w},y_{l})$ the DPO loss can be written as (Yan et al., 2025):
@@ 3.2 Failure Mode of DPO when $\pi_{\theta}(y\mid x)$ is Close to Zero
$\ell_{\mathrm{DPO}}=\log\!\Bigl(1+\bigl[\alpha\,z^{\beta}\bigr]\Bigr),\qquad\alpha=\Bigl[\tfrac{\pi_{ref}(y_{w}\!\mid x)}{\pi_{ref}(y_{w}\!\mid x)}\Bigr]^{\beta},\;z=\tfrac{\pi_{\theta}(y_{l}\!\mid x)}{\pi_{\theta}(y_{l}\!\mid x)}.$
Denoting $\pi^{+}\!=\!\pi_{\theta}(y_{w}\!\mid x)$ and $\pi^{-}\!=\!\pi_{\theta}(y_{l}\!\mid x)$, differentiation gives
@@
Hence, we have that $\lvert\partial_{\pi^{-}}\ell\rvert/\lvert\partial_{\pi^{+}}\ell\rvert=\pi^{+}/\pi^{-}$. The gradient on the rejected probability therefore grows much faster than that on the chosen probability; as $\pi^{-}\!\to 0$ we have $\partial_{\pi^{-}}\ell\!\to\!\infty$ but $\partial_{\pi^{+}}\ell\!\to 0$. Because chosen and rejected completions usually share many tokens, the vanishing positive gradient cannot offset the dominant negative gradient, so the model is driven to lower both probabilities. This synchronous likelihood collapse—observed empirically in late-stage DPO training—degrades overall performance (Pal et al., 2024).
3.3 Stability of the MIO Gradient
@@
If the DV/MINE mutual information estimator causes gradient starvation when $\pi_{\theta}(y\mid x)$ is small, replacing it with a JSD-based mutual information estimator eliminates this issue. In other words, when $\pi_{\theta}(y\mid x)$ approaches zero, MIO does not suffer from vanishing policy gradients. This claim can be verified directly by differentiating the MIO loss. The MIO objective for the same triple is
@@
with $\operatorname{LR}^{\pm}=\log\pi_{\theta}^{\pm}-\log\pi_{ref}^{\pm}$. Writing $\sigma^{\pm}=\sigma\!\bigl(\beta\operatorname{LR}^{\pm}\bigr)$, one obtains
@@
4 Experiment
4.1 DPO Breaks When $\pi_{\theta}(y_{l}\mid x)$ near 0, But MIO Not!
The previous section theoretically demonstrated that methods dependent on $I_{\mathrm{DV}}(\theta)$, such as DPO and PPO, encounter issues with policy gradients becoming ineffective when $\pi_{\theta}(y\mid x)$ approaches zero, especially for $\pi_{\theta}(y_{l}\mid x)$. This is particularly evident because $\pi_{\theta}(y_{l}\mid x)$ is more likely to approach zero during training than $\pi_{\theta}(y_{w}\mid x)$. However, this problem does not occur in the MIO approach. The earlier discussion was purely theoretical, and while it is challenging to validate this in practice, we have prepared a toy model experiment to empirically investigate this issue.
Following (Yan et al., 2025), the toy model setup, shown in Figure 1, consists of a discrete space with 4 prompts and 10 responses. The policy $\pi_{\theta}$, implemented as a three-layer MLP, processes a one-hot input vector and outputs a categorical distribution over the responses, which are divided into three categories: selected responses (first 4 dimensions), rejected responses (next 4 dimensions), and unseen responses (final 2 dimensions). Each prompt has a corresponding optimal response (e.g., response 1 is optimal for prompt 1). In DPO or MIO training, preference data is generated through a mini-batch sampling strategy, where an ideal annotator matches each prompt with its optimal response, creating a diagonal preference matrix. In each mini-batch, one additional response is randomly selected to form preference pairs, ensuring diverse gradient updates.
Figure 1: Toy model setup(Yan et al., 2025). Left: the optimal policy where the highlighted blocks represent optimal responses. Right: preference dataset construction.
Figure 2 contrasts the learning dynamics of DPO with those of our proposed MIO objective. When the probabilities of both the chosen and rejected responses remain in a moderate range, DPO converges smoothly and the likelihood of the chosen response declines only marginally relative to that of the rejected one. Once the rejected–response probability becomes very small, however, DPO develops a pathological synchronous collapse: the probabilities of both the chosen and the rejected responses decrease in lock-step. Across all regimes we tested, MIO displays no such joint degradation. When the rejected probability is pushed towards zero, DPO and its derivatives exhibit synchronous decline of the chosen and rejected likelihoods, mirroring what is observed in practice  (Yan et al., 2025; Chen et al., 2024; Ho & Ermon, 2016), whereas MIO remains entirely immune to this failure mode. A formal explanation of this qualitative difference is provided in Section 3.2.
The toy model serves as an abstract simulation that amplifies the 3D-properties of DPO (Yan et al., 2025), which leads to a decrease in the likelihood of both chosen and rejected responses. However, MIO does not exhibit this effect. While the toy model setup differs significantly from real-world LLM training—such as in sampling frequency—it provides valuable insights. In real-world scenarios, DPO and MIO are typically trained over one epoch, with each data point used only a few times. In contrast, the toy model samples the same data points repeatedly. Conceptually, this is akin to treating each input/output as a token rather than a full prompt/response, where each token may be sampled multiple times during real-world training.
4.2 Real World Experimental Setup
We fine-tune both Mistral-7B-SFT(Tunstall et al., 2023a) and LLaMA3-8B-SFT on the 64k-prompt UltraFeedback-Binarized (Cui et al., 2024) corpus, as well as train directly on Qwen2.5-7B-Instruct, where each prompt is paired with a chosen and a rejected completion selected by GPT-4, yielding 64k positive–negative pairs; we retain the official train/validation split of UltraFeedback-Binarized.
Model quality is assessed on eight public mathematical and reasoning suites: Hendrycks MATH (Hendrycks et al., 2021a) (12.5 k competition-level problems plus the level-5 “hard” subset), Minerva Math (Lewkowycz et al., 2022a) (quantitative-reasoning items distilled from GSM8K and MATH), MultiMedQA (Singhal et al., 2022) (six medical QA benchmarks spanning USMLE, PubMedQA, and consumer health queries), MathQA (Amini et al., 2019a) (37 k multiple-choice word problems with executable rationales), GSM8K (Cobbe et al., 2021a) (8.5 k grade-school arithmetic questions), AQuA-RAT (Ling et al., 2017) (100 k algebraic problems with free-text solutions), and MuSR (Sprague et al., 2024) (narrative tasks requiring multistep soft reasoning over 700–1000-word stories). The code, hyperparameter settings, and evaluation metrics is provid in the Appendix  G.
Table 1: Evaluation results on 8 benchmarks. The best and second best are marked.
[…]
Mistral-7B-SFT
SFT 0.14 8.80 49.03 29.91 32.29 19.29 2.72 39.94
DIL(Xiao et al., 2025) 0.04 6.46 49.54 28.94 28.35 21.33 2.11 38.36
DPO(Rafailov et al., 2023) 0.04 8.14 49.21 30.72 36.47 19.69 2.72 43.65
IPO(Azar et al., 2023) 0.08 6.50 48.39 29.41 3.71 19.69 0.23 35.71
KTO(Ethayarajh et al., 2024) 0.04 10.38 49.76 30.56 41.75 21.65 3.43 43.39
NCA(Chen et al., 2024) 0.14 8.92 49.58 30.59 37.68 18.50 2.49 39.68
ORPO(Hong et al., 2024) 0.04 9.68 49.40 29.78 36.39 24.80 2.57 42.06
SimPO(Meng et al., 2024) 0.19 8.36 49.16 30.93 39.20 18.11 2.49 39.02
SLiC(Zhao et al., 2023) 0.08 8.72 49.74 30.66 39.73 20.47 3.17 42.99
CPOXu et al. (2024a) 0.16 7.14 49.18 27.57 31.59 23.37 3.40 42.44
MIO 0.21 9.08 49.82 30.96 40.16 24.83 3.57 43.42
LLaMA3-8B-SFT
SFT 0.06 14.06 56.84 28.07 50.26 25.90 4.61 39.94
DIL(Xiao et al., 2025) 0.19 17.16 57.01 34.85 55.86 23.62 5.06 40.75
DPO(Rafailov et al., 2023) 0.04 16.48 55.54 28.17 54.38 23.62 4.23 39.11
IPO(Azar et al., 2023) 0.12 15.26 56.99 29.85 53.07 25.69 4.91 36.77
KTO(Ethayarajh et al., 2024) 0.04 16.42 31.36 19.83 55.57 22.05 5.03 35.05
NCA(Chen et al., 2024) 0.08 17.24 57.53 28.41 54.81 25.20 4.46 38.62
ORPO(Hong et al., 2024) 0.08 16.86 53.81 28.01 48.98 25.59 4.68 38.76
SimPO(Meng et al., 2024) 0.06 17.16 57.13 36.35 53.37 24.80 4.61 41.08
SLiC(Zhao et al., 2023) 0.17 15.18 51.94 28.58 46.22 26.77 3.59 38.03
CPOXu et al. (2024a) 0.13 14.26 54.01 26.13 52.16 25.39 3.85 39.15
MIO 0.20 17.68 57.32 36.85 56.86 27.62 5.44 42.81
Qwen-7B-Instruct
Instruct 0.04 25.20 55.54 33.33 71.19 21.65 49.47 38.5
DIL(Xiao et al., 2025) 0.12 23.12 55.31 33.66 72.21 24.80 46.22 37.29
DPO(Rafailov et al., 2023) 0.11 26.16 57.36 33.21 73.46 22.83 50.08 39.55
IPO(Azar et al., 2023) 0.18 25.44 56.28 34.20 71.57 21.26 50.38 40.08
KTO(Ethayarajh et al., 2024) 0.19 26.91 52.86 29.7 75.08 20.08 46.00 37.99
NCA(Chen et al., 2024) 0.07 25.34 56.61 34.30 72.56 22.05 49.47 39.81
ORPO(Hong et al., 2024) 0.02 26.34 55.49 37.98 74.30 28.58 32.18 39.97
SimPO(Meng et al., 2024) 0.19 26.6 45.25 34.80 71.72 25.98 38.44 37.96
SLiC(Zhao et al., 2023) 0.06 27.54 57.32 36.31 73.99 25.98 46.07 38.76
CPOXu et al. (2024a) 0.02 18.76 54.76 36.14 72.71 28.74 46.27 37.17
MIO 0.22 27.84 57.61 38.17 75.39 30.16 51.36 40.34
The result is in table 1. For Mistral-7B-SFT, MIO achieves the best performance on five tasks and ranks second on two others across the eight mathematical and reasoning benchmarks. For LLaMA3-8B-SFT, MIO ranks first in seven of the eight benchmarks and second in one. For Qwen2.5-7B-Instruct, MIO achieves the best performance across all eight benchmarks. These results demonstrate that MIO offers superior generalization in reasoning and mathematical problem-solving compared to existing alignment methods.
4.3 MIO prevents the chosen-reward collapse
Figure 3: Training curves on Mistral-7B-SFT. Unlike DPO, IPO, and ORPO, MIO increases the rewards of chosen responses while suppressing rejected ones, thereby avoiding the “synchronous collapse” observed in prior methods.
During DPO training, the rewards of all responses ($y_{w}$ and $y_{l}$) often decrease simultaneously, a harmful “unlearning” effect that degrades downstream performance (Xiao et al., 2025; Meng et al., 2024; Pal et al., 2024). Figure 3 shows that the proposed MIO objective completely eliminates this failure mode: the chosen-rewards increase while the rejected rewards drop slightly. Section 3.2 provides a theoretical explanation of this behavior.
Some readers may argue that NCA-Pair can also increase the chosen reward while decreasing the rejected reward. To examine this, Figure  5 (in Appendix) presents a direct comparison between MIO and NCA-Pair in terms of reward margin and related metrics. The results show that MIO achieves a greater increase in both the chosen reward and the reward margin, as well as a larger decrease in the rejected reward, compared to NCA-Pair. This indicates that MIO leverages the reward signal more effectively than NCA-Pair.
@@ 5 Conclusion
In this work, we propose a novel perspective that frames RLHF and DPO as a unified form of contrastive learning over the distributions of chosen and rejected responses, grounded in mutual information. Building on this connection, we provide rigorous explanations for several phenomena observed in prior work. For example, reasoning paths that are not sampled by the base model remain inaccessible even after RL training, and DPO fails to update effectively when the probability of a rejected response becomes too small. We show that all of these issues can be attributed to the DV/MINE mutual information estimator. Furthermore, we introduce the Mutual Information Optimization (MIO), replacing the Donsker-Varadhan (DV) bound with a Jensen-Shannon (JSD) bound. MIO demonstrates superior preservation of reasoning and mathematical abilities and offers greater training stability compared to existing baselines. Extensive experimental results demonstrate that MIO achieves superior performance on a wide range of mathematical and reasoning benchmarks. We hope that our work could offer useful insights for the community and help address the limitations of current RLHF methods. This work does not explore the effects of alternative mutual information estimators, which we leave as a promising direction for future research. Details regarding LLM usage are provided in the appendix J.
Ethics Statement
This work adheres to the ICLR Code of Ethics. No human-subject or animal experiments were conducted. All datasets used, including UltraFeedbackBinarized (Cui et al., 2023a; Tunstall et al., 2023b), were obtained and used in compliance with their licenses and applicable policies. We took precautions to avoid privacy risks and discriminatory outcomes. No personally identifiable information was processed, and no experiments posed security or safety concerns. We are committed to transparency and integrity throughout the research lifecycle.
Reproducibility Statement
We took extensive steps to ensure the results are reproducible. We release code, configuration files, and processing scripts in an anonymized repository to support replication. The experimental setup—including training procedures, model architectures, hyperparameters, and hardware details—is documented in the paper and the repository README. We also provide a detailed description of experiment details G to facilitate faithful reimplementation. Furthermore, we have released all code associated with training and experiments, which is publicly available at https://anonymous.4open.science/r/MIO-63E6/ for full reproducibility.
@@
Substitute equation A.3 into equation A.1.
@@
Insert equation A.4 into equation A.2.
@@
and using equation A.3 in equation A.5 we arrive at
@@
For every prompt $x$ draw $y_{w}\sim\pi_{\text{chosen}}(\cdot\!\mid\!x)$ (positive) and $y_{l}\sim\pi_{\text{rejection}}(\cdot\!\mid\!x)$ (negative). Fix a critic $T_{\phi}\!\in\!\mathcal{F}$. With $z^{+}=T_{\phi}(x,y_{w})$ and $z^{-}=T_{\phi}(x,y_{l})$, define the two DV / InfoNCE estimators:
@@
Write the logit gap $\Delta\coloneqq z^{+}-z^{-}$. With $\sigma(u)=\tfrac{1}{1+e^{-u}}$ we have $\hat{I}^{+}=\log\sigma(\Delta),\;\hat{I}^{-}=\log\sigma(-\Delta)$.
@@
Because $\sigma(u)>0\;\forall u\in\mathbb{R}$,
@@
$\nabla_{\theta}\hat{I}^{+}=\frac{\partial\hat{I}^{+}}{\partial\Delta}\,\nabla_{\theta}\Delta,\qquad\nabla_{\theta}\hat{I}^{-}=\frac{\partial\hat{I}^{-}}{\partial\Delta}\,\nabla_{\theta}\Delta.$
Combine equation B.3–equation B:
@@
The scalar factor $\sigma(\Delta)/\sigma(-\Delta)\,<0$; hence
$\boxed{\;\left\langle\nabla_{\theta}\hat{I}^{+},\,\nabla_{\theta}\hat{I}^{-}\right\rangle<0\;}$
i.e. the two gradient vectors always point in opposite directions.
Consequence (first‑order).
$\hat{I}^{+}(\theta_{\text{new}})>\hat{I}^{+}(\theta)\;\Longrightarrow\;\hat{I}^{-}(\theta_{\text{new}})<\hat{I}^{-}(\theta).$
Second‑order coupling: extrema coincide
Shared critical points.
From equation B, $\nabla_{\theta}\hat{I}^{+}=0$ iff $\nabla_{\theta}\hat{I}^{-}=0$; thus both objectives share the same set of stationary points.
Hessian analysis.
Let $f(\theta)=\hat{I}^{+}(\theta)=h(\Delta(\theta))$ and $g(\theta)=\hat{I}^{-}(\theta)=h(-\Delta(\theta))$ with $h(u)=\log\sigma(u)$. Then
$h^{\prime}(u)=\sigma(-u)>0,\qquad h^{\prime\prime}(u)=-\sigma(u)\sigma(-u)<0.$
Using standard composition rules,
$\nabla^{2}_{\theta}f$ $=h^{\prime\prime}(\Delta)\,\nabla\Delta\nabla\Delta^{\!\top}+h^{\prime}(\Delta)\,\nabla^{2}\!\Delta,$ (B.9)
$\nabla^{2}_{\theta}g$ $=h^{\prime\prime}(-\Delta)\,\nabla\Delta\nabla\Delta^{\!\top}-h^{\prime}(-\Delta)\,\nabla^{2}\!\Delta.$ (B.10)
At any common critical point $\theta^{\star}$ we have $\nabla\Delta(\theta^{\star})=0$; the rank‑one terms vanish, leaving
$\nabla^{2}_{\theta}f(\theta^{\star})=h^{\prime}(\Delta^{\star})\,\nabla^{2}\!\Delta(\theta^{\star}),\qquad\nabla^{2}_{\theta}g(\theta^{\star})=-h^{\prime}(-\Delta^{\star})\,\nabla^{2}\!\Delta(\theta^{\star}).$
Because $h^{\prime}(\cdot)>0$, the two Hessians differ only by an overall sign.
Extremum equivalence.
If $\nabla^{2}_{\theta}f(\theta^{\star})\prec 0$ (negative definite), then $\nabla^{2}_{\theta}g(\theta^{\star})\succ 0$ (positive definite); hence
@@
Implication for optimisation.
Therefore,
$\arg\max_{\theta}\,\hat{I}^{+}(\theta)\;=\;\arg\min_{\theta}\,\hat{I}^{-}(\theta),$
provided the optimum is strict and isolated. Gradient‑based methods that converge to a strict maximiser of $\hat{I}^{+}$ will, by construction, converge to the corresponding strict minimiser of $\hat{I}^{-}$, rigorously validating the coupling intuition expressed in the main text.
Appendix C Mutual-Information Estimation: Practical Surrogate and Constrain Critic T